OpenStudy (anonymous):

Solve the following equation. Then place the correct number in the box provided. 2x < 6

OpenStudy (anonymous):

divide both sides by 2

OpenStudy (anonymous):

divide both side with 2

OpenStudy (anonymous):

to get x < 3

OpenStudy (anonymous):

so, x<3

OpenStudy (anonymous):

so then 3 will be my answer?

OpenStudy (anonymous):

\[ \frac{ 2x }{ 2 } <\frac{ 6 }{ 2 }\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it depends . but in actual case you answer will be numbers from \[-\infty \to 2.99\]

OpenStudy (anonymous):

it can be written as \[(-\infty, 3)\]

OpenStudy (anonymous):

oh ok, thank you all so much! Can ya'll help me with a few more please?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can i have a medal?

OpenStudy (anonymous):

Solve the following inequality. Then place the correct answer in the box provided. Answer in terms of an mixed number. 8P + 2 > 3P - 15

OpenStudy (anonymous):

collect like terms

OpenStudy (anonymous):

open another question